OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (anonymous):

There are two other answers.

OpenStudy (anonymous):

y=1/bx+b(a-1) y=-1/bx+b(1-a)

OpenStudy (anonymous):

y = −bx − b(1 − a) y = bx − b(a − 1)

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{general formula}\hspace{.33em}\\~\\ y-y_1=m(x-x_1) \hspace{.33em}\\~\\ \normalsize \text{equation of line passing through points (a,b)}\hspace{.33em}\\~\\ \normalsize \text{and having a slope of b is}\hspace{.33em}\\~\\ y-b=b(x-a) \hspace{.33em}\\~\\ \normalsize \text{try to match it with ur options or simplify }\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (arabpride):

Don't forget to close this post if you're done. Thanks! :3

OpenStudy (anonymous):

I'm not done, I still don't get it. @arabpride

OpenStudy (anonymous):

what do you not get @peytonamberxoxo

OpenStudy (nincompoop):

To solve for the equation of a line when a slope, m, is provided and a point (x1, y1), use the point-slope formula y-y1= m(x-x1) and then solve for y

OpenStudy (nincompoop):

In your problem, Slope= b Point (a,b) Use the point-slope form (y-b) = b(x-a) Use distributive property in the right-hand side of the equation: (y-b) = bx - ba Then solve for y to get the equation the line using the slope-intercept form y = bx - ba + b